# On the Brun–Titchmarsh theorem. I

### v3

@@ 1.3. Connections with Landau–Siegel zeros
holds with some constant $\delta>0$, the real zero $\beta$ (if exists) of $L(s,\chi)$ then satisfies
@@
On the other hand, it is believed that $\pi(x,a;q)\leqslant 2\mathrm{Li}(x)/\varphi(q)$ as long as $q$ is at most a small power of $x$; see [Mo12] for the details and history. Recently, Maynard [Ma13] demonstrated this expectation is valid for all $q^{8}\leqslant x.$ His method involves a log-free zero density estimate for Dirichlet $L$-functions and careful analysis of Landau–Siegel zeros. Up to a very small constant $\varepsilon$, we have shown in Theorem 1.5 that $\pi(x,a;q)$ is at most twice as much as the conjectural bound, provided that $q<x^{5/12},$ $q$ is squarefree and all prime factors are sufficiently small.
@@
Acknowledgements
We are very grateful to Étienne Fouvry and Kaisa Matomäki for their kind comments and suggestions. This work is supported in part by NSFC (No. 12025106).
@@ 2. Dirichlet polynomials and character sums
For an arbitrary coefficient $\boldsymbol{\alpha}=(\alpha_{n})_{n\sim N},$ we have
@@
For an arbitrary coefficient $\boldsymbol{\alpha}=(\alpha_{n})_{n\sim N}$ and $V>0,$ we have
#$\#\Big{\{}\chi\,(\operatorname{mod}{q}):\Big{|}\sum_{n\sim N}\alpha_{n}\chi(n)\Big{|}>V\Big{\}}\ll\|\boldsymbol{\alpha}\|^{2}NV^{-2}+\|\boldsymbol{\alpha}\|^{6}q^{1+\varepsilon}NV^{-6}.$
@@
Let $F\in\mathcal{C}_{0}^{\infty}(\mathbf{R}^{2})$ be a Schwartz function with support in $]L,2L]\times]L,2L]$, satisfying
@@
Let $\eta,q,N,L\geqslant 1.$ Suppose $-β\boldsymbol{\alpha},\bbeta$ are complex coefficients supported in $[N,2N],$ and $F$ is a Schwartz function defined as above. Then
@@
×$\qquad\times((M+N)^{\frac{1}{4}}+(KMN)^{\frac{1}{40}}(KM+KN)^{\frac{1}{8}}).$
@@
If the coefficients $-β\boldsymbol{\alpha},\bbeta$ have suitable convolution structures, we can do a bit better than Theorem 2.1.
@@
for any $\varepsilon,A>0,$ with an implied constant depending only on $(\varepsilon,A,c)$. Here $\vartheta\leqslant 7/64$ is the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$
@@
At first glance, the restriction that $\boldsymbol{\alpha}^{\prime\prime}$ and $-β′′\bbeta^{\prime\prime}$ vanish on the set $\{n\geqslant 2:\omega(n)>(\log N)^{1/3}\}$ is quite artificial. In our applications (as well as many other situations) this is harmless since only a minority of integers have been ruled out. We remark that this condition is designed for the later application of a trick due to Fouvry (see Lemma 7.1 below).
The control on the error term in Theorem 2.2 depends heavily on well-factorizations of $-β\boldsymbol{\alpha},\bbeta.$ The proof will be given in Section 7, and we appeal to Kloostermania in the following version.
@@
Let $C,M,N\geqslant 1,$ $a,r,s\in\mathbf{Z}^{+}$ pairwise coprime with $\mu(rs)\neq 0,$ and $g(c,m,n)$ a smooth function with compact support in $[C,2C]\times[M,2M]\times[N,2N]$ such that
@@
For an arbitrary complex coefficient $\boldsymbol{\beta}=(\beta_{n}),$ we have
@@
$\Delta$ $=a^{\vartheta}(rsCM)^{\frac{1}{2}}\frac{(1+1/X)^{\vartheta}}{1+X^{2}}\Big{(}1+X+\frac{C}{rM}\Big{)}^{\frac{1}{2}}\Big{(}1+X+\frac{N}{rs}\Big{)}^{\frac{1}{2}}+s^{-1}MN^{\frac{1}{2}}$
@@
Theorems 2.1 and 2.2 allow us to derive the following versions of Lindelöf Hypothesis on average.
@@
Let $N,L,q\geqslant 3$ and $g$ a smooth function with compact support in $[1,2],$ satisfying $\|g^{(j)}\|_{\infty}\ll\eta^{j}$ for some $\eta\geqslant 1.$ Let $\boldsymbol{\lambda}=(\lambda_{n})$ be an arbitrary coefficient with support in $[1,N].$ Suppose that $N,L\ll q^{O(1)}.$ Then we have
@@
Let $N_{1},N_{2},N,L,q\geqslant 3$ and $g$ a smooth function with compact support in $[1,2],$ satisfying $\|g^{(j)}\|_{\infty}\ll\eta^{j}$ for some $\eta\geqslant 1.$ Let $\boldsymbol{\lambda}=\boldsymbol{\lambda}_{1}*\boldsymbol{\lambda}_{2}$ be a convolution, where $\boldsymbol{\lambda}_{1},$ $\boldsymbol{\lambda}_{2}$ have supports in $[1,N_{1}],$ $[1,N_{2}]$ with $N_{1}N_{2}=N$ and $\|\boldsymbol{\lambda}_{1}\|_{\infty},\|\boldsymbol{\lambda}_{2}\|_{\infty}\leqslant 1.$ We also assume that $\boldsymbol{\lambda}$ vanishes on the set $\{n\geqslant 2:\omega(n)>(\log N)^{1/3}\}.$ Suppose that $N,L\ll q^{O(1)}.$ Then we have
$\frac{1}{\varphi(q)}$ $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{n}\frac{\lambda_{n}\chi(n)}{\sqrt{n}}\Big{|}^{2}\Big{|}\sum_{l\in\mathbf{Z}}g\Big{(}\frac{l}{L}\Big{)}\frac{\chi(l)}{\sqrt{l}}\Big{|}^{2}$
$\ll\eta^{\frac{9}{2}}q^{\frac{\vartheta-1}{2}}N_{1}^{\frac{1}{2}}N_{2}^{2}(N_{1}q^{-\frac{1}{2}}+1)q^{\varepsilon}+\eta q^{\varepsilon}+\eta^{\frac{9}{2}}N_{1}q^{-\frac{1}{2}+\varepsilon}$
for any $\varepsilon,A>0,$ where the implied constant depends only on $(\varepsilon,A).$
@@
Let $q$ be a large prime and $N\geqslant 1$. Let $\boldsymbol{\alpha}=(\alpha_{n})_{n\leqslant N}$ be an arbitrary coefficient supported on squarefree numbers and $\|\boldsymbol{\alpha}\|_{\infty}\leqslant 1.$ Then we have
@@ 3.1. Trace functions over finite fields
For an odd prime $p$, we consider the middle-extension sheaves on $\mathbf{A}^{1}_{\mathbf{F}_{p}}$. Recall that a sheaf $\mathcal{F}$ on $\mathbf{A}^{1}_{\mathbf{F}_{p}}$ is said to be middle-extension, if $\mathcal{F}\simeq j_{*}(j^{*}\mathcal{F})$ holds for each $j:U\rightarrow\mathbf{A}^{1}$ as the open immersion of a non-empty open set on which $\mathcal{F}$ is lisse. Fix an isomorphism $\iota:\overline{\mathbf{Q}}_{\ell}\rightarrow\mathbf{C}$ with another prime $\ell\neq p$. Following Katz [Ka88, Section 7.3.7], we define trace functions as follows.
@@
where $\mathrm{Frob}_{x}$ denotes the geometric Frobenius at $x\in\mathbf{F}_{p},$ and $V_{\mathcal{F}}$ is a finite dimensional $\overline{\mathbf{Q}}_{\ell}$-vector space, which is corresponding to a continuous finite-dimensional Galois representation and unramified at every closed point $x$ of $U.$
@@
where $S(\mathcal{F})\subset\mathbf{P}^{1}(\overline{\mathbf{F}}_{p})$ denotes the ((((finite)))) set of singularities of $\mathcal{F}$, and $\mathrm{Swan}_{x}(\mathcal{F})$ $(\geqslant 0)$ denotes the Swan conductor of $\mathcal{F}$ at $x$.
@@
  • Let $f\in\mathbf{F}_{p}(X)$ be a rational function, and $\psi$ a primitive additive character on $\mathbf{F}_{p}$, then $\psi(f(x))$ is a trace function of an $\ell$-adic middle-extension sheaf, which is taken to be zero when meeting a pole of $f$ at $x$. More precisely, one can show that there exists an $\ell$-adic middle-extension sheaf modulo $p$, denoted by $\mathcal{L}_{\psi(f)},$ such that $x\mapsto\psi(f(x))$ is the trace function of $\mathcal{L}_{\psi(f)}.$ The conductor can be bounded in terms of the degree of $f$, independent of $p$.
@@
    According to Deligne, there exists an $\ell$-adic middle-extension sheaf $\mathcal{K}\ell$, called a Kloosterman sheaf, such that $\mathrm{Kl}(\cdot,p)$ gives its trace function over $\mathbf{F}_{p}^{\times}.$ The sheaf $\mathcal{K}\ell$ is geometrically irreducible and is of rank $2$, the conductor of which is bounded by $5$.
An middle-extension sheaf on $\mathbf{P}_{\mathbf{F}_{p}}^{1}$, which is pointwise pure of weight 0 is said to be a Fourier sheaf if no geometrically irreducible component is geometrically isomorphic to an Artin–Schreier sheaf $\mathcal{L}_{\psi}$ attached to some additive character $\psi$ of $\mathbf{F}_{p}.$
@@
  • The Fourier transform is involutive, in the sense that we have a canonical arithmetic isomorphism $\mathrm{FT}_{\psi}(\mathcal{G})\simeq[\times(-1)]^{*}\mathcal{F},$ where [×$[\times(-1)]^{*}$ denotes the pull-back by the map $x\mapsto-x;$
@@
Let $\mathcal{K}\ell$ be the Kloosterman sheaf of rank $2$ as above. For all $a,b\in\mathbf{F}_{p}^{\times}$ with $a\neq b,$ the tensor product :=([×$\mathcal{F}:=([\times a]^{*}\mathcal{K}\ell)\otimes([\times b]^{*}\mathcal{K}\ell)$ is a Fourier sheaf, which is geometrically irreducible, pure of weight zero, of rank $4,$ and has the geometric monodromy group $\mathrm{SL}_{2}\times\mathrm{SL}_{2}.$ The conductor of $\mathcal{F}$ is at most $25$
Moreover, for each non-trivial additive character $\psi$ of $\mathbf{F}_{p},$ the Fourier transform $\mathrm{FT}_{\psi}(\mathcal{F})$ is also Fourier sheaf, which is geometrically irreducible and of conductor at most $6250$
@@
Before going into more details on trace functions over finite fields, we would like mention another definition of Kloosterman sums defined over $\mathbf{Z}/q\mathbf{Z}:$ for $q\geqslant 1$, put
@@
where $m,n\in\mathbf{Z}$ and $a\overline{a}\equiv 1\,(\operatorname{mod}{q}).$ It is clear that $\mathrm{Kl}(x,p)\sqrt{p}=S(x,1;p)=S(1,x;p)$ for each prime $p$, if we identify $\mathbf{F}_{p}$ by $\{0,1,\cdots,p-1\}.$ We say $S(*,0;q)$ a Ramanujan sum, which satisfies $|S(m,0;q)|\leqslant(m,q);$ in general, we have the Weil bound
@@
Let $M,N\geqslant 1$ and $V_{1},V_{2}$ two fixed smooth functions with compact supports in $[1,2].$ We then have
@@
$S(\mathbf{n})=\sum_{x\,(\operatorname{mod}{p})}\prod_{1\leqslant j\leqslant 2\nu}\mathrm{Kl}(n_{j}x).$
@@
We also expect there exist squareroot cancellations if the coordinates of $\mathbf{n}$ appear in suitable configurations. More precisely, according to [FGKM14] or [Xi17], we have
@@ 3.2. Composite trace functions and arithmetic exponent pairs
By the Chinese remainder theorem, a trace function can be also defined for squarefree moduli, which allows Wu, Xi and Sawin [WX21] to develop the general form of $q$-analogue of van der Corput method and arithmetic exponent pairs.
@@
where $K_{p}$ is the trace function of certain $\ell$-adic middle-extension sheaf $\mathcal{F}_{p}$ on $\mathbf{A}^{1}_{\mathbf{F}_{p}}$ for each $p\mid q$. We adopt the convention that $K_{q}(n)=1$ for all $n$ if $q=1.$ In practice, the value of $K_{p}(n)$ may depend on the complementary divisor $q/p.$ Originally, $K_{p}$ is a function living in $\mathbf{Z}/p\mathbf{Z}$; one can extend its domain to $\mathbf{Z}$ by periodicity, and one can do the same thing to $K_{q}$.
@@
Roughly speaking, the Pólya–Vinogradov method can be used to bound $\mathfrak{S}(K,W)$ non-trivially for $|I|>q^{\frac{1}{2}+\varepsilon}$, and it was observed by Heath-Brown [HB78] that one may go further if $q$ has food factorizations, and this led him to prove a Weyl-type subconvexity bound for $L(\tfrac{1}{2},\chi)$ to smooth moduli. Inspired by Heath-Brown, Wu and Xi [WX21] developed the method of arithmetic exponent pairs for averages of composite trace functions that may go far beyond the Pólya–Vinogradov bound, as long as the factorization of $q$ is good enough. We now summarize results of the arithmetic exponent pairs developed in [WX21]. For more backgrounds and applications of the $q$-analogue of van der Corput method and arithmetic exponent pairs, the readers may refer to [WX21] directly and references therein.
@@
    it is pointwise pure of weight 00;
@@
  • $I$ is an interval with $|I|<q\delta;$
@@
[…]
@@
[…]

### v4

@@ 1.3. Connections with Landau–Siegel zeros
holds with some constant $\delta>0$, the real zero $\beta$ (if it exists) of $L(s,\chi)$ then satisfies
@@
On the other hand, it is believed that $\pi(x;q,a)\leqslant 2\mathrm{Li}(x)/\varphi(q)$ as long as $q$ is at most a small power of $x$; see [Mo12] for the details and history. Recently, Maynard [Ma13] demonstrated this expectation is valid for all $q^{8}\leqslant x.$ His method involves a log-free zero density estimate for Dirichlet $L$-functions and careful analysis of Landau–Siegel zeros. Up to a very small constant $\varepsilon$, we have shown in Theorem 1.5 that $\pi(x;q,a)$ is at most twice the conjectural bound, provided that $x^{5/12}, qq$ is squarefree and all prime factors are sufficiently small.
@@ 2. Dirichlet polynomials and character sums
For any complex sequence $\boldsymbol{\alpha}=(\alpha_{n})_{n\sim N},$ we have
@@
For any complex sequence $\boldsymbol{\alpha}=(\alpha_{n})_{n\sim N}$ and $V>0,$ we have
$\#\Big\{\chi\,(\operatorname{mod}{q}):\Big|\sum_{n\sim N}\alpha_{n}\chi(n)\Big|>V\Big\}\ll\|\boldsymbol{\alpha}\|^{2}NV^{-2}+\|\boldsymbol{\alpha}\|^{6}q^{1+\varepsilon}NV^{-6}.$
@@
Let $F\in\mathcal{C}_{0}^{\infty}(\mathbf{R}^{2})$ be a Schwartz function with support in $L,2L]\times]L,2L]$, satisfying
@@
[…]
@@
$\qquad\times((M+N)^{\frac{1}{4}}+(KMN)^{\frac{1}{40}}(KM+KN)^{\frac{1}{8}}).$
@@
[…]
@@
for any $\varepsilon>0,$ with an implied constant depending only on $(\varepsilon,c)$. Here $\vartheta\leqslant 7/64$ is the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$
@@
[…]
@@
Let $C,M,N\geqslant 1,$ $a,r,s\in\mathbf{Z}^{+}$ pairwise coprime with $\mu(rs)\neq 0,$ and $g(c,m,n)$ a smooth function with compact support in $C,2C]\times[M,2M]\times[N,2N]$ such that
@@
For any complex sequence $\boldsymbol{\beta}=(\beta_{n}),$ we have
@@
$\Delta$ $=a^{\vartheta}(rsCM)^{\frac{1}{2}}\frac{(1+1/X)^{\vartheta}}{1+X^{\frac{1}{2}}}\Big(1+X+\frac{C}{rM}\Big)^{\frac{1}{2}}\Big(1+X+\frac{N}{rs}\Big)^{\frac{1}{2}}+s^{-1}MN^{\frac{1}{2}}$
@@
Theorems 2.1 and 2.2 allow us to derive the following versions of Lindelöf Hypothesis (or square-root cancellation) on average.
@@
Let $N,L,q\geqslant 3$ and let $g$ be a smooth function with compact support in $[1,2],$ satisfying $\|g^{(j)}\|_{\infty}\ll\eta^{j}$ for some $\eta\geqslant 1.$ Let $\boldsymbol{\lambda}=(\lambda_{n})$ be a complex sequence with support in $[1,N].$ Suppose that $N,L\ll q^{O(1)}.$ Then we have
@@
[…]
$\frac{1}{\varphi(q)}$ $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big|\sum_{m}\frac{\alpha_{m}\chi(m)}{\sqrt{m}}\Big|^{2}\Big|\sum_{n}\frac{\beta_{n}\chi(n)}{\sqrt{n}}\Big|^{2}\Big|\sum_{l\in\mathbf{Z}}g\Big(\frac{l}{L}\Big)\frac{\chi(l)}{\sqrt{l}}\Big|^{2}$
$\ll\eta^{\frac{9}{2}}(q^{\frac{\vartheta-1}{2}}M^{\frac{1}{2}}N^{2}+1)(Mq^{-\frac{1}{2}}+1)q^{\varepsilon}$
One may rewrite the product of two Dirichlet polynomials in Lemma 2.6 as one Dirichlet polynomial with convoluted coefficients. This special convolution structure allows us to obtain square-root cancellations on average for the Dirichlet polynomial longer than Lemma 2.5. In this sense, Lemma 2.6 is new. We will give more precise comparisons in Section 8.4. See also [BPZ20] for the second moment of Dirichlet $L$-functions twisted by Dirichlet polynomials, but with asymptotic formulae.
@@
Let $q$ be a large prime and $N\geqslant 1$. Let $\boldsymbol{\alpha}=(\alpha_{n})_{n\leqslant N}$ be a complex sequence supported on squarefree numbers and $\|\boldsymbol{\alpha}\|_{\infty}\leqslant 1.$ Then we have
@@ 3.1. Trace functions over finite fields
For an odd prime $p$, we consider middle-extension sheaves on $\mathbf{A}^{1}_{\mathbf{F}_{p}}$. Recall that a sheaf $\mathcal{F}$ on $\mathbf{A}^{1}_{\mathbf{F}_{p}}$ is said to be a middle-extension sheaf if there exists a non-empty open subset $U\subset\mathbf{A}^{1}_{\mathbf{F}_{p}}$ and an open immersion $j:U\hookrightarrow\mathbf{A}^{1}_{\mathbf{F}_{p}}$ such that $\mathcal{F}$ is lisse on $U$ and $\mathcal{F}\simeq j_{*}(j^{*}\mathcal{F})$. Fix an isomorphism $\iota:\overline{\mathbf{Q}}_{\ell}\rightarrow\mathbf{C}$ for a $\ell\neq p$. Following Katz [Ka88, Section 7.3.7], we define trace functions as follows.
@@
where $\mathrm{Frob}_{x}$ denotes the geometric Frobenius at $x\in\mathbf{F}_{p},$ and $V_{\mathcal{F}}$ is a finite dimensional $\overline{\mathbf{Q}}_{\ell}$-vector space, which corresponds to a continuous finite dimensional Galois representation and unramified at every closed point $x$ of $U.$
@@
where $S(\mathcal{F})\subset\mathbf{P}^{1}(\overline{\mathbf{F}}_{p})$ denotes the ((finite)) set of singularities of $\mathcal{F}$, and $\mathrm{Swan}_{x}(\mathcal{F})$ $(\geqslant 0)$ denotes the Swan conductor of $\mathcal{F}$ at $x$.
@@
  • Let $f\in\mathbf{F}_{p}(X)$ be a rational function, and $\psi$ a primitive additive character on $\mathbf{F}_{p}$, then $\psi(f(x))$ is a trace function of an $\ell$-adic middle-extension sheaf, which is taken to be zero when $x$ is a pole of $f$. More precisely, one can show that there exists an $\ell$-adic middle-extension sheaf modulo $p$, denoted by $\mathcal{L}_{\psi(f)},$ such that $x\mapsto\psi(f(x))$ is the trace function of $\mathcal{L}_{\psi(f)}.$ The conductor can be bounded in terms of the degree of $f$, independent of $p$.
@@
    According to Deligne, there exists an $\ell$-adic middle-extension sheaf $\mathcal{K}\ell$, called a Kloosterman sheaf, such that $\mathrm{Kl}(\cdot,p)$ gives its trace function over $\mathbf{F}_{p}^{\times}.$ The sheaf $\mathcal{K}\ell$ is geometrically irreducible of rank $2$, with conductor bounded by $5$.
A middle-extension sheaf on $\mathbf{P}_{\mathbf{F}_{p}}^{1}$ that is pointwise pure of weight 0 is said to be a Fourier sheaf if no geometrically irreducible component is geometrically isomorphic to an Artin–Schreier sheaf $\mathcal{L}_{\psi}$ attached to some additive character $\psi$ of $\mathbf{F}_{p}.$
@@
  • The Fourier transform is involutive, in the sense that we have a canonical arithmetic isomorphism $\mathrm{FT}_{\psi}(\mathcal{G})\simeq[\times(-1)]^{*}\mathcal{F},$ where $\times(-1)]^{*}$ denotes the pull-back by the map $x\mapsto-x;$
@@
Let $\mathcal{K}\ell$ be the Kloosterman sheaf of rank $2$ as above. For all $a,b\in\mathbf{F}_{p}^{\times}$ with $a\neq b,$ the tensor product $\mathcal{F}:=([\times a]^{*}\mathcal{K}\ell)\otimes([\times b]^{*}\mathcal{K}\ell)$ is a Fourier sheaf, which is geometrically irreducible, pure of weight zero, of rank $4,$ and has the geometric monodromy group $\mathrm{SL}_{2}\times\mathrm{SL}_{2}.$ The conductor of $\mathcal{F}$ is at most $25.$
Moreover, for each non-trivial additive character $\psi$ of $\mathbf{F}_{p},$ the Fourier transform $\mathrm{FT}_{\psi}(\mathcal{F})$ is also a Fourier sheaf, which is geometrically irreducible and of conductor at most $6250.$
@@
Before going into more details on trace functions over finite fields, we would like to mention another definition of Kloosterman sums defined over $\mathbf{Z}/q\mathbf{Z}:$ for $q\geqslant 1$, put
@@
where $m,n\in\mathbf{Z}$ and $a\overline{a}\equiv 1\,(\operatorname{mod}{q}).$ It is clear that $\mathrm{Kl}(x,p)\sqrt{p}=S(x,1;p)=S(1,x;p)$ for each prime $p$, if we identify $\mathbf{F}_{p}$ with $\{0,1,\cdots,p-1\}.$ We call S(,0;q)S(*,0;q) a Ramanujan sum, which satisfies $|S(m,0;q)|\leqslant(m,q);$ in general, we have the Weil bound
@@
Let $M,N\geqslant 1$, and let $V_{1},V_{2}$ be two fixed smooth functions with compact supports in $[1,2].$ We then have
@@
$S(\mathbf{n})=\sum_{x\,(\operatorname{mod}{p})}\prod_{1\leqslant j\leqslant 2\nu}\mathrm{Kl}(n_{j}x,p).$
@@
We also expect there exist square-root cancellations if the coordinates of $\mathbf{n}$ appear in suitable configurations. More precisely, according to [FGKM14] or [Xi17], we have
@@ 3.2. Composite trace functions and arithmetic exponent pairs
By the Chinese remainder theorem, a trace function can also be defined for squarefree moduli, which allows Wu, Xi and Sawin [WX21] to develop the general form of $q$-analogue of van der Corput method and arithmetic exponent pairs.
@@
where $K_{p}$ is the trace function of a certain $\ell$-adic middle-extension sheaf $\mathcal{F}_{p}$ on $\mathbf{A}^{1}_{\mathbf{F}_{p}}$ for each $p\mid q$. We adopt the convention that $K_{q}(n)=1$ for all $n$ if $q=1.$ In practice, the value of $K_{p}(n)$ may depend on the complementary divisor $q/p.$ Originally, $K_{p}$ is a function living in $\mathbf{Z}/p\mathbf{Z}$; one can extend its domain to $\mathbf{Z}$ by periodicity, and one can do the same thing to $K_{q}$.
@@
Roughly speaking, the Pólya–Vinogradov method can be used to bound such sums non-trivially for $|I|>q^{\frac{1}{2}+\varepsilon}$, and it was observed by Heath-Brown [HB78] that one may go further if $q$ has good factorizations, and this led him to prove a Weyl-type subconvexity bound for $L(\tfrac{1}{2},\chi)$ to smooth moduli. Inspired by Heath-Brown, Wu–Xi–Sawin [WX21] developed the method of arithmetic exponent pairs for averages of composite trace functions that may go far beyond the Pólya–Vinogradov bound, as long as the factorization of $q$ is good enough. We now summarize results of the arithmetic exponent pairs developed in [WX21]. For more background and applications of the $q$-analogue of van der Corput method and arithmetic exponent pairs, the readers may refer to [WX21] directly and references therein.
@@
    it is pointwise pure of weight 0;
@@
  • II is an interval with $N:=|I|<q\delta;$
@@
(3.6) $0\leqslant\kappa\leqslant\frac{1}{2}\leqslant\lambda\leqslant 1,\ \ 0\leqslant\nu\leqslant 1$
@@
Keep the notation and convention as above. For each compositely universally amiable trace function $K_{q},$ the estimate (3.2) holds for all